OpenStudy (anonymous):

Fred can paint a house in 10 hours. Together, Pat and Fred can paint the house in 8 hours. Which equation, when solved, will produce the number of hours needed for Pat to paint the house alone? Choose one answer. a. 10/8+9/8=1 b. 8/10+8/p=1 c. 8/10+9/8=1 d. 10/8+8/p=1 Please show work too! and is it B? Not quite sure

ganeshie8 (ganeshie8):

you're right !

ganeshie8 (ganeshie8):

F paints the house in 10 hours => F paints 1/10 th of the house in 1 hour both P + F paints in 8 hours => P + F paint 1/8 th of the house in 1 hour => 1/10 + 1/P = 1/8 => 8/10 + 8/P = 1

OpenStudy (anonymous):

ok, thanks :)

ganeshie8 (ganeshie8):

np :)